OpenStudy (anonymous):

Will I actually get an answer on this one? Or will I wait for nothing like last question! A 5 mile cab ride cost 7.20 a 9 mile cab ride cost 11.60 find a linear equation that models a relationship between cost c and distance d A. c=1.44d+4.40 B. d=1.10c+4.40 C. c=1.29d+1.70 D. c=1.10d+1.70

OpenStudy (kropot72):

The distance d is equivalent to the x variable, and the cost c is equivalent to the y variable. So we can find the slope from the ordered pairs (5, 7.2) and (9, 11.6). Are you with me?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

\[m=\frac{11.6-7.2}{9-5}=you\ can\ calculate\]

OpenStudy (anonymous):

2.4

OpenStudy (kropot72):

Not really. The numerator simplifies to 4.4 and the denominator simplifies to 4. What did you get?

OpenStudy (kropot72):

\[m=\frac{11.6-7.2}{9-5}=\frac{4.4}{4}=?\]

OpenStudy (anonymous):

1.4

OpenStudy (anonymous):

its a correct

OpenStudy (kropot72):

No. m = 4.4/4 = 1.1 Now we have a value for m, the linear equation can be written as: c = 1.10d + b where b is a constant that needs to be found by substition as follows: 7.2 = (1.10 * 5) + b ........(1) Can you solve equation (1) tnd the value of b ?

OpenStudy (kropot72):

substitution*

OpenStudy (anonymous):

I got d

OpenStudy (kropot72):

Well you can find that d is the correct choice without finding the value of the constant. But you will learn more by finding the value of b.

OpenStudy (anonymous):

We have two points (5, 7.20) and (9, 11.60) So use the slope formula to get m = (y2-y1)/(x2-x1) m = ( 11.60- 7.20)/(9-5) m = (4.4)/4 m = 1.1 So the slope of the line that passes through the two points is 1.1